# Prediction of Cellular Malignancy Using Electrical Impedance Signatures and Supervised Machine Learning

### 05 Mar 2026 v3

@@ Abstract
Bioelectrical properties of cells such as relative permittivity, conductivity, and characteristic time constants vary significantly between healthy and malignant cells across different frequencies. These distinctions provide a promising foundation for diagnostic and classification applications. This study systematically reviewed 20 scholarly articles to compile 535 datasets of quantitative bioelectric parameters and evaluated their utility in predictive modeling. Three supervised machine learning algorithms- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) were implemented and tuned using key hyperparameters to assess classification performance. Model effectiveness was evaluated using accuracy and F1 score as performance metrics. Results demonstrate that Random Forest achieved the highest predictive accuracy of   90% when configured with a maximum depth of 4 and 100 estimators. These findings highlight the potential of integrating bioelectrical property analysis with machine learning for improved diagnostic decision-making. Similarly, for KNN and SVM, the F1 score peaked at approximately 78% and 76.5%, respectively. Future work will explore incorporating additional discriminative features, leveraging stimulated datasets, and optimizing hyperparameter through advanced search strategies. Ultimately, hardware prototype with embedded micro-electrodes and real-time control systems could pave the path for practical diagnostic tools capable of in-situ cell classification.
@@ 1 Introduction
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. When combined with equivalent circuit model and electromagnetic analysis, these measurements can be transformed into physically meaningful dielectric and conductive parameter.By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [Biau2016],[Rigatti2017],[Guo2003],[Hossain2024Authorea],[Suthaharan2016]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods. Fig.1 shows the schematic of the sequence of steps for the proposed model discussed earlier. Bioelectrical measurements -such as impedance, conductivity, and dielectric loss -are acquired from test samples using microelectrodes. These features are then input into a pretrained machine learning classifier for binary classification, distinguishing between malignant and healthy cells based on their electrical signatures.
Figure 1: Schematic representation of the proposed models, in which bioelectrical measurements are obtained from test samples and then processed using pre-trained machine learning (ML) models.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Fig. 2(d–f) illustrates the corresponding electric field magnitude distribution inside the homogeneous spheres calculated using equation (3):
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [Giannoukos2014]. Fig. 3(a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig.3 (b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
@@
The Cole-Cole representation therefore provides a physically grounded framework for comparing dielectric signatures for malignant and non-tumorigenic breast cell lines, enabling visualization of dispersion characteristics and subsequent quantitative classification analysis. The Cole-Cole relaxation model provides a robust framework for characterizing dielectric dispersion in biological tissues [Cole1941]. Fig. 4 illustrates the complex permittivity distribution of malignant (T-47D, MCF-7, MDA-MB-231) and non-tumorigenic MCF-10A breast epithelial cell lines, alongside reported tissue baselines [Ref82, Ref83, Ref84]. The theoretical Cole-Cole relaxation loci (gray arcs) are superimposed to conceptualize the dispersion behavior.
@@ 4 Results and Discussion
Fig. 5 (a) –(i) presents the quantitative evaluation results, including (i) class-wise and macro-averaged performance metrics, (ii) the corresponding confusion matrices, and (iii) training and cross validation learning curves for three representative hyperparameter configurations: Case I : Maximum depth=5, Number of estimators =100; Case II: Maximum depth=3, Number of estimators=100, and Case III : Maximum depth=15, number of estimators =500. In all three cases, the training accuracy consistently exceeds the cross-validation accuracy, indicating the presence of moderate overfitting. However, the performance gap decreases for deeper trees, suggesting improved generalization with increased model capacity. A comparative assessment of Fig. 5 (a), (d) and (g) reveals that the RF classifier demonstrates strong discriminative capacity for malignant cells (Class III), achieving F1 scores exceeding 0.88 in Case III. Conversely, the classification performance for benign cells (Class II) remains comparatively lower, with F1 scores of approximately 0.59 and 0.54 for Case I and II, respectively and improving to approximately 0.71 in Case III. The relatively lower benign -class sensitivity highlights the inherent difficulty in distinguishing benign from malignant cellular dielectric signatures. Enhancing benign-class discrimination is clinically significant as improved early stage differentiation may facilitate timely diagnosis and optimized therapeutic interventions.
@@
K-Nearest Neighbor (KNN) 78% $k=2$, number of iterations = 100
@@ 5 Conclusion and Future Work
Bioelectrical differences in cell properties such as relative permittivity, conductivity, and characteristic time constant as a function of frequency demonstrated significant variation between malignant and healthy cells. These distinctions highlight the potential of this technology for applications in cell classification and diagnosis. In this study, thirty-three scholarly articles were reviewed, and their datasets were compiled to provide a comprehensive overview and quantitative analysis of cell properties. Three supervised ML models- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) model were evaluated for their effectiveness in predicting and classifying cell types. Model hyperparameters including :(i) maximum depth and number of estimators for RF, (ii) kernel type and regularization parameters for SVM, and (iii) number of neighbors for KNN, were tuned to improve the predictive performance. Accuracy and F1 scores served as evaluation metrics to compare model effectiveness. Among the models, RF , an ensemble-based model achieved the highest performance with an accuracy of approximately 90% at a maximum depth of 4 and 100 estimators. Future research should focus on expanding the parameter space by integrating additional discriminative bioelectrical features and leveraging datasets derived from both experimental and simulation studies. Advanced hyperparameter optimization techniques, such as grid search and Bayesian methods, could further enhance predictive performance. Moreover, the development of hardware prototypes with embedded microelectrode arrays and real-time control systems may enable point-of-care applications for rapid and reliable diagnosis.

### 13 Mar 2026 v4

@@ Abstract
Bioelectrical properties of cells such as relative permittivity, conductivity, and characteristic time constants vary significantly between healthy and malignant cells across different frequencies. These distinctions provide a promising foundation for diagnostic and classification applications. This study systematically reviewed 20 scholarly articles to compile 535 datasets of quantitative bioelectric parameters in the kHz-MHz frequency range and evaluated their utility in predictive modeling. Three supervised machine learning algorithms- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) were implemented and tuned using key hyperparameters to assess classification performance. In the second stage, a physics informed framework was incorporated to derive additional dielectric descriptors such as imaginary permittivity, loss tangent and charge relaxation time from the measured parameters. Random Forest based feature importance analysis was employed to identify the most discriminative dielectric parameters influencing the classification process. The results indicate that dielectric loss related parameters, particularly imaginary permittivity and conductivity, contribute significantly to the classification of cellular states. While the incorporation of physics-derived features improves model interpretability and reduces overfitting tendencies, the overall classification accuracy remains comparable to models trained using primary dielectric descriptors. The proposed approach highlights the potential of physics-informed machine learning for improving the analysis of dielectric spectroscopy data in the biomedical diagnostics.
@@ 1 Introduction
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. When combined with equivalent circuit model and electromagnetic analysis, these measurements can be transformed into physically meaningful dielectric and conductive parameter.By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [Biau2016],[Rigatti2017],[Guo2003],[Hossain2024Authorea],[Suthaharan2016]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods.
In this work, a comparative investigation is conducted to evaluate the performance of these machine learning (ML) algorithms for the classification of biological cell types using dielectric spectroscopy data measured across the kHz-MHz frequency range. In the first stage, the ML models are trained using primary dielectric properties; namely relative permittivity and electrical conductivity to assess their baseline classification capability based on experimentally measured parameters.
In the second stage, a physics-informed framework is incorporated to derive additional dielectric descriptors including: the charge relaxation time ($\tau$), loss tangent ($\tan\delta$), and imaginary permittivity ($\varepsilon^{\prime\prime}$). These parameters are associated with dielectric relaxation phenomena and energy dissipation mechanisms within biological tissues. The influence of these physics-derived features on the classification performance of the ML models is systematically investigated under identical hyperparameter configuration. Furthermore, a feature importance analysis based on the Random Forest algorithm is performed to identify the most discriminative dielectric parameters contributing to accurate cell classification. The proposed analysis therefore provides insight into the role of physics-informed dielectric parameters in improving machine-learning-based characterization of biological cells.
The overall workflow of the proposed physics-informed framework is illustrated in Fig. 1.
Figure 1: Overall workflow of the proposed methodology for dielectric spectroscopy-based cell classification. (a) Initial machine learning study using experimentally reported dielectric parameters compiled from multiple studies, including frequency, relative permittivity, and electrical conductivity. Supervised learning algorithms such as Random Forest, Support Vector machine , and K-Nearest Neighbor are used for classification and evaluation. (b) Physics informed feature engineering framework where additional dielectric descriptors such as imaginary permittivity, loss tangent and charge relaxation time are derived from the measured parameters. Feature importance analysis is then applied to identify the influential descriptors before retaining the machine learning model for improved classification.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Fig.  2 (d–f) illustrates the corresponding electric field magnitude distribution inside the homogeneous spheres calculated using equation (3):
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [Giannoukos2014]. Fig.  3 (a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig. 3 (b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
@@
The Cole-Cole representation therefore provides a physically grounded framework for comparing dielectric signatures for malignant and non-tumorigenic breast cell lines, enabling visualization of dispersion characteristics and subsequent quantitative classification analysis. The Cole-Cole relaxation model provides a robust framework for characterizing dielectric dispersion in biological tissues [Cole1941]. Fig.  4 illustrates the complex permittivity distribution of malignant (T-47D, MCF-7, MDA-MB-231) and non-tumorigenic MCF-10A breast epithelial cell lines, alongside reported tissue baselines [Ref82, Ref83, Ref84]. The theoretical Cole-Cole relaxation loci (gray arcs) are superimposed to conceptualize the dispersion behavior.
@@ 2.5 Understanding K-Nearest Neighbor
2.6 Physics Based Machine Learning Approach
In the second part of the study, the original dataset is augmented using a physics-based modeling framework to derive additional dielectric descriptors from the measured parameters. Specifically, secondary features such as (i) loss tangent (ii) imaginary permittivity and (iii) charge relaxation time are computed from the primary dielectric properties, namely electrical conductivity and relative permittivity of the biological cell samples. The objective of incorporating these physics-derived parameters is to enrich the feature space and provide additional physical insight into dielectric relaxation and energy dissipation mechanisms within the cells, thereby improving the robustness of the machine learning model when limited primary descriptors are available.
@@ 3 Data Collection and Methodology
In the second stage of this study, physics-based model framework is implemented using the same datasets. The complete feature set is constructed by compiling six parameters: (i) measured frequency, (ii) electrical conductivity, (iii) relative permittivity, (iv) loss tangent, (v) imaginary permittivity, and (vi) relaxation time. A Random Forest classifier is then employed to evaluate the relative contribution of each parameter to the classification task. The model is trained using 500 estimators with a maximum tree depth of 15, and the corresponding feature importance weights are compared to determining the significance of each dielectric parameter in accurately predicting the cancer state.
Subsequently, a feature selection strategy is applied by imposing an importance threshold greater than 0.1 to retain only the most discriminative parameters. The reduced feature set is then used to retain the Random Forest model using the same hyper parameter configuration. The resulting model performance is evaluated using element wise classification metrics and validation testing curves to assess the improvement in prediction efficiency achieved through the incorporation of physics-informed feature selection.
@@ 4 Results and Discussion
Fig.  5 presents the quantitative evaluation results, including (i) class-wise and macro-averaged performance metrics, (ii) the corresponding confusion matrices, and (iii) training and cross validation learning curves for three representative hyperparameter configurations: Case I : Maximum depth=5, Number of estimators =100; Case II: Maximum depth=3, Number of estimators=100, and Case III : Maximum depth=15, number of estimators =500. In all three cases, the training accuracy consistently exceeds the cross-validation accuracy, indicating the presence of moderate overfitting. However, the performance gap decreases for deeper trees, suggesting improved generalization with increased model capacity. A comparative assessment of Fig. 5 (a), (d) and (g) reveals that the RF classifier demonstrates strong discriminative capacity for malignant cells (Class III), achieving F1 scores exceeding 0.88 in Case III. Conversely, the classification performance for benign cells (Class II) remains comparatively lower, with F1 scores of approximately 0.59 and 0.54 for Case I and II, respectively and improving to approximately 0.71 in Case III. The relatively lower benign -class sensitivity highlights the inherent difficulty in distinguishing benign from malignant cellular dielectric signatures. Enhancing benign-class discrimination is clinically significant as improved early stage differentiation may facilitate timely diagnosis and optimized therapeutic interventions.
@@
Fig.  7 presents the performance evaluation of the K-Nearest Neighbor (KNN) classifier for different values of the neighborhood parameters, k. Subfigures (a), (c), and (e) illustrate the corresponding classification reports showing class-wise precision, recall, and F1 score, while subfigures (b), (d) and (f) present the associated confusion matrices. The classification is performed across all three cell categories: Class I (normal), Class II (benign), and Class III (malignant).
The results indicate that the classifier performs comparatively well for the malignant cell class (Class III), achieving an F1 score of approximately 0.88 when k=2. This improved performance can be attributed to the relatively large number of samples available for this class in the dataset. However, as the number of neighbors increases to k=10 and k=15, the classification performance for the benign class (Class II) deteriorates significantly, as reflected by the reduction in precision and recall values.
Figure 7: Performance evaluation of the K-Nearest Neighbor (KNN) classifier for different neighborhood sizes. (a), (c) and (e) show the classification reports indicating precision, recall, and F1 score for each class, while (b), (d) and (f) present the corresponding confusion matrices. The models are evaluated for k=2, k=10, and k=15, respectively
@@
K-Nearest Neighbor (KNN) 79% $k=2$, number of iterations = 100
@@
Fig. 8 presents the evaluation results obtained after incorporating additional physics-derived dielectric parameters into the dataset. Fig. 8(a) illustrates the relative feature importance obtained using the Random Forest Classifier. The height of each bar represents the contribution of the corresponding feature towards the classification task. From the feature importance analysis, it can be observed that the imaginary component of permittivity ($\varepsilon^{\prime\prime}$), electrical conductivity ($\sigma$), and loss tangent ($\tan\delta$) exhibit high importance scores, indicating that dielectric loss-related parameters play a significant role in distinguishing between different cell types.
Figure 8: (a)Feature importance scores obtained using Random Forest classifier for the extended dielectric feature set. (b) Feature selection algorithm used to retain parameters above the importance threshold (c) Confusion matrix obtained for the optimized Random Forest model with 500 estimators and maximum depth of 15. (d) Corresponding classification report showing precision, recall and F1 score for each cell class.
Initially the classification model was trained using the complete set of features including: (i) measured frequency, (ii) conductivity, (iii) relative permittivity, (iv) loss tangent, (v) imaginary permittivity, and (vi) relaxation time ($\tau$). The feature selection process retained parameters with an importance score of greater than 0.005, as illustrated in the algorithm shown in Fig. 8 (b). The Random Forest Classifier was implemented using 500 estimates and a maximum depth of 15.
The corresponding classification performance is summarized in Fig. 8(c) and (d) , which present the confusion matrix of element-wise evaluation metrics, respectively. The model achieved an overall classification accuracy of approximately 82%. For malignant cell classification (Class III) the model obtained an F1 score of 90%, precision of 97% and recall of 84%. Although the overall accuracy is slightly lower than the results obtained from the previous Random Forest model using the same hyperparameter, a notable improvement is observed in the recall score for the benign cell class, which increased from 75% to 93%. The recall metric is particularly important in biomedical classification problems, as it measures the ability of the machine learning model to correctly identify all relevant instances of a given class. An improvement in recall therefore indicates enhanced sensitivity of the model in detecting the corresponding cell state, which is critical for early diagnosis applications.
Figure 9: Evaluation of the physics informed Random Forest model with reduced model complexity (estimators=100, maximum depth=5) : (a) Classification report for the full feature set; (b) Learning curve showing training and validation scores; (c) Feature importance ranking of the dielectric parameters; and (d) Classification performance using only the selected high-importance physics derived parameters.
To further evaluate the influence of physics-informed feature selection and model complexity on the classification performance, an additional analysis was performed using a reduced Random Forest architecture with the number of estimates set to 100 and maximum tree depth limited to 5 as shown in Fig.  9. The corresponding feature importance weights for all dielectric parameters are summarized in Fig.  9(c). Based on a feature importance threshold of 0.1, three dominant parameters were identified (i) imaginary permittivity (= 0.19), (ii) electrical conductivity (=0.20) and (iii) relative permittivity (=0.18).
Fig.  9 (a) presents the classification performance obtained when the complete feature set was used in the Random Forest model. The model achieved an overall accuracy of approximately 77%, which represents a marginal improvement compared to the earlier model accuracy of 76%. However, analysis of the learning curves shown in Fig.  9(b) indicates a reduction in overfitting behavior, as evidenced by the smaller gap between the training and validation scores. Subsequently, the model was retained using only the three dominant physics-informed parameters identified through the feature importance analysis. The resulting model achieved an overall accuracy of approximately 75%, indicating only a slight reduction in predictive performance. Furthermore, no significant improvement was observed in the corresponding F1 scores for the individual classes.
These results suggest that while physics-derived dielectric parameters contribute valuable interpretability and reduce model complexity, their inclusion does not necessarily lead to significant improvements in classification accuracy for the current dataset.
@@ 5 Conclusion and Future Work
Bioelectrical differences in cell properties such as relative permittivity, conductivity, and characteristic time constant as a function of frequency demonstrated significant variation between malignant and healthy cells. These distinctions highlight the potential of this technology for applications in cell classification and diagnosis. In this study, twenty scholarly articles were reviewed, and their datasets were compiled to provide a comprehensive overview and quantitative analysis of cell properties. Three supervised ML models- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) model were evaluated for their effectiveness in predicting and classifying cell types. Model hyperparameters including :(i) maximum depth and number of estimators for RF, (ii) kernel type and regularization parameters for SVM, and (iii) number of neighbors for KNN, were tuned to improve the predictive performance. Accuracy and F1 scores served as evaluation metrics to compare model effectiveness. The second part of the study investigated the potential of integrating physics-derived dielectric descriptors with machine learning algorithms for the classification of biological cell states. The results demonstrate that dielectric loss-related parameters such as imaginary permittivity, conductivity, and loss tangent exhibit strong discriminative capability for distinguishing between normal, benign, and malignant cells. Although the inclusion of physics-informed features improves model interpretability and reduces overfitting tendencies, the overall classification accuracy remains comparable to models trained solely on primary dielectric parameters. These findings highlight the importance of combining physics-based insights with machine learning approaches for informed understanding of dielectric spectroscopy- based biomedical diagnostics.
Future research should focus on expanding the parameter space by integrating additional discriminative bioelectrical features and leveraging datasets derived from both experimental and simulation studies. Advanced hyperparameter optimization techniques, such as grid search and Bayesian methods, could further enhance predictive performance. Moreover, the development of hardware prototypes with embedded microelectrode arrays and real-time control systems may enable point-of-care applications for rapid and reliable diagnosis.